OpenStudy (anonymous):

Given: ∆BCA is a right triangle. Prove: a2 + b2 = c2 The two-column proof with missing justifications proves the Pythagorean Theorem using similar triangles. Statement Justification Draw an altitude from point C to Let = a = b = c = h = x = y y + x = c a2 = cy; b2 = cx a2 + b2 = cy + b2 a2 + b2 = cy + cx a2 + b2 = c(y + x) a2 + b2 = c(c) a2 + b2 = c2 Which is NOT a justification for the proof? Addition Property of Equality Pythagorean Theorem Pieces of Right Triangles Similarity Theorem Cross Product Property

OpenStudy (anonymous):

Statement Justification Draw an altitude from point C to Let segment BC = a segment CA= b segment AB= c segment CD= h segment DB= x segmet AD= y y + x = c c over a = a over y , c over b = b over x a2 = cy; b2 = cx a2 + b2 = cy + b2 a2 + b2 = cy + cx a2 + b2 = c(y + x) a2 + b2 = c(c) a2 + b2 = c2

OpenStudy (anonymous):

^^^^^^^^^^^ better version of second half of the question

OpenStudy (anonymous):

@SmoothMath ?

OpenStudy (anonymous):

@robtobey ?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

which one do you think it is

OpenStudy (anonymous):

pythagorean was my guess, to be honest all of them seemed like they could be the answer, it has me very confused

jimthompson5910 (jim_thompson5910):

you are correct, you're trying to prove the pythagorean theorem so there's no way you can use something you're trying to prove as justification (that would be using circular reasoning)

OpenStudy (anonymous):

okay thank you very much (:

jimthompson5910 (jim_thompson5910):

np